OpenStudy (anonymous):

Use the substitution method to solve the given system of equations. 4x - 8y =8 x- 5= 2y a.(0,1) b.infinitely many solutions c.no solution d.(5,0)

OpenStudy (anonymous):

is the answer c

OpenStudy (campbell_st):

4x - 8y = 8 (1) x - 5 = 2y (2) rearrange equation (1) 4(2y) = 4(x - 4) or 2y = x - 4 substitute into (2) x - 5 = x - 4 -5 = -4 the system of equations has no solutions this is because the lines are parallel both have a gradient of 1/2 and intercepts if y = -5/2 (2) and y = -1 (1)

OpenStudy (anonymous):

thanks :)